An MSU linebacker of mass 114.0 kg sacks a UM quarterback of mass 95.0 kg. Just after they collide, they are momentarily stuck together, and both are moving at a speed of 2.60 m/s. If the quarterback was at rest just before he was sacked, how fast was the linebacker moving just before the collision?
Answer #1

momentum conservation

Pi = Pf

=> 114*u1 + 95*u2 = 114*v1 + 95*v2

=> 114*u1 + 95*0 = 114+95)*2.60

=> u1 = 4.76 m/s , fast was the linebacker moving just before the collision